OpenStudy (alex_mattucci):

Which expression is equivalent to the area of metal sheet required to make this square-shaped traffic sign? One side of the sign is 4x+3. a. 16x2 + 12x + 9 b. 16x2 + 24x + 9 c. 8x2 + 9 d. 8x2 + 6

OpenStudy (anonymous):

what do you know about the sides of a square?

OpenStudy (alex_mattucci):

what do u mean?

OpenStudy (anonymous):

just what i asked... you know what a sqaure is, right? what can you tell me about its sides?

OpenStudy (alex_mattucci):

it has four sides, each are equal in length

OpenStudy (anonymous):

excellent! and what about the area of a square?

OpenStudy (alex_mattucci):

length times width

OpenStudy (alex_mattucci):

ohhhhh

OpenStudy (anonymous):

but for a square, length and width are...?

OpenStudy (alex_mattucci):

they are the same

OpenStudy (anonymous):

man, you are on it!!!

OpenStudy (alex_mattucci):

so it would be (4x+3)(4x+3)

OpenStudy (anonymous):

excellent!!! and what do you get when you multiply that out?

OpenStudy (alex_mattucci):

c. 8x2 + 9

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

do you know the distributive property?

OpenStudy (alex_mattucci):

sorry, i coppied and pasted the wrong one from above lol...

OpenStudy (anonymous):

a(b+c) = ab + ac

OpenStudy (alex_mattucci):

The right one is b. 16x2 + 24x + 9

OpenStudy (anonymous):

however, if a is a binomial (two terms) also, it goes like this... suppose a = f+g then we get ab + ac = (f+g)b + (f+g) c and using distribution again we get... fb + gb + fc + gc

OpenStudy (anonymous):

there you go... you know how to multiply those out, right? you're not just guessing...

OpenStudy (alex_mattucci):

yeah i do

OpenStudy (anonymous):

awesome! hey, just a hint... whenever you can draw a picture to represent what is going on in a problem, do so. make it sufficicently large and put all available info on it. we humans are visual creatures!

OpenStudy (alex_mattucci):

alright

OpenStudy (alex_mattucci):

ty so much

OpenStudy (anonymous):

good luck and have a great day!